OpenStudy (anonymous):

A cashier offers Nathan an opportunity to save twenty-five percent on his purchase. The original cost is eighty-three dollars. Estimate what his savings would be, then calculate the actual savings.

OpenStudy (stamp):

25% of $83 = savings Estimate. $83 ~ $80 25% of $83 ~ 25% of 80 = 20 Find actual savings. 25% of $83 = 0.25 of 83 0.25 of 83 = 0.25 (83) 0.25 (83) = 83/4 Simplify. 83/4 = 20 3/4 20 3/4 = 20.75 $20.75 saved